Grade 10Algebra

If zebra were younger by 5years than what she really is then square of her age would have been 11more than 5times her actual age .what is her age now

Vinay
7 Years agoGrade 10

1 Answer

Anantha vishnu R
7 Years ago
Let  the actual age be x then from the question
(x – 5)2 = 11 + 5x
ie; x2 – 10x + 25 = 11+ 5x
By solving this we get x = 11 or x = 1.